SAT verbal scores are normally distributed with a mean 433 and a standard deviation of 90. Use the Empirical Rule to determine what percent of scores lies between 433 and 523.

Answer by stanbon(75887)

SAT verbal scores are normally distributed with a mean 433 and a standard deviation of 90. Use the Empirical Rule to determine what percent of scores lies between 433 and 523.
---------
z(433) = (433-433)/90 = 0
--
z(523) = (523-433)/90 = 90/90 = 1
----
P(433< x < 523) = P(0< z < 1) = 34%
----------------
Cheers,
Stan H.
---------------